OpenStudy (anonymous):

length of curve y = e^x from 0 to 1

OpenStudy (anonymous):

Do it the same as before

OpenStudy (jamesj):

Use the same formula we just used in your last problem and see how far you can get. Ask when you're stuck, but I think it's fair that we don't work this problem for you from start to finish.

OpenStudy (anonymous):

I'm at \[\int\limits_{0}^{1}\sqrt{1 + e ^{2x}}\] and I'm stuck

OpenStudy (jamesj):

Use the most banal substitution you can imagine: u = sqrt(1 + e^(2x) ) and work this through patiently.

OpenStudy (anonymous):

how would i take the integral of it

OpenStudy (jamesj):

I remember this integral very well: it was the last problem in my HS textbook. It's not hard, but you need to go through several steps.

OpenStudy (anonymous):

find du/dx

OpenStudy (anonymous):

okay so u = \[\sqrt{1 + e ^{2x}} \] so du = \[e ^{2x}/\sqrt{1+e ^{2x}}\]

OpenStudy (jamesj):

You dropped a 2

OpenStudy (anonymous):

but the 1/2 and 2 cancel out

OpenStudy (jamesj):

my bad

OpenStudy (anonymous):

okay but where do i go from there?

OpenStudy (jamesj):

You want an integral in u, so express that derivative in u

OpenStudy (anonymous):

but there is no du in the equation

OpenStudy (anonymous):

Remember an integral looks like Sf(x)dx

OpenStudy (jamesj):

We're using a substition

OpenStudy (anonymous):

yeah but i set u as the whole integral, so du isn't in the integral

OpenStudy (anonymous):

no, u is only sqrt(1 + e^(2x) )

OpenStudy (anonymous):

there is still a dx hanging there...

OpenStudy (anonymous):

which is the whole equation

OpenStudy (anonymous):

okay so how do i put du back in the equation?

OpenStudy (anonymous):

yeah i know how to do u substitution i just don't understand how to do it int his problem

OpenStudy (jamesj):

We've used the simpliest, dumbest substitution possible, f(x) = \sqrt(1 + e^(2x) ) = u, so g(u) = u

OpenStudy (jamesj):

So what now is du/dx in terms of u? You need to calculate this and I really think if it takes you 5 minutes to figure it out, good, do it.

OpenStudy (jamesj):

I know ti took me a while the first time as well.

OpenStudy (anonymous):

what if i set u = e^x then did \[\int\limits_{0}^{1}\sqrt{1 + u ^{2}} \] which is just arctan u

OpenStudy (jamesj):

No, it is NOT arctan u.

OpenStudy (jamesj):

Use the substitution we gave you. It works.

myininaya (myininaya):

i did a trig substitution and now i'm using partial fractions lol it does seem to be long

OpenStudy (jamesj):

Yes it is long.

OpenStudy (anonymous):

okay so right now i have integral u, how do i put in a du?

OpenStudy (jamesj):

You need to find an expression for du/dx in terms of u and then use the substitution rule for integrals I wrote above.

OpenStudy (jamesj):

So try!

OpenStudy (anonymous):

You have almost got it, mate

OpenStudy (jamesj):

I've got to go; conference call.

OpenStudy (anonymous):

Have fun conference calling

OpenStudy (anonymous):

i just don't understand how to put a du in, i'd have to put in the whole derivative somehow

OpenStudy (anonymous):

Just substitute it in James' formula

OpenStudy (anonymous):

I.e., integral of f(x) dx = u . dx/du du So actually you need an expression for dx/du in terms of u; just the reciprocal of du/dx

myininaya (myininaya):

i think james's way is easiest so he said let u=(something)^(1/2) then u^2=something now taking derivative of both sides we get...?

OpenStudy (jamesj):

Here's an example problem, so you can the method. Evaluate \[I = \int\limits \sqrt{4 + e^x} \ dx\] Let \[u = \sqrt{4 + e^x}\] then \[e^x = u^2 - 4\] and \[\frac{du}{dx} = \frac{e^x}{2 \sqrt{4+e^x}} = \frac{u^2 - 4}{2u}\] Thus \[I = \int\limits u(x) dx = \int\limits u(x) \frac{dx}{du} du = \int\limits \frac{2u^2}{u^2 - 4} du\]

OpenStudy (jamesj):

Now, \[\frac{2u^2}{u^2 - 4} = 2 + \frac{8}{u^2 - 4} = 2 + 2 \left( \frac{1}{u- 2} - \frac{1}{u+2} \right)\] Thus \[I = 2u + \ln(u-2) - \ln(u+2) + C\]

OpenStudy (jamesj):

Corrections: I dropped a couple of 2s \[I = 2u + 2\ln(u - 2) - 2 \ln(u + 2)\]

OpenStudy (jamesj):

Now substitute back in u(x) = sqrt(4 + e^x)

myininaya (myininaya):

\[I = \int\limits\limits \sqrt{4 + e^x} \ dx \] \[u=\sqrt{4+e^x} => u^2=4+e^x =>2 u du=e^x dx\] \[I=\int\limits_{}^{} \frac{e^x}{e^x} \sqrt{4+e^x} dx=\int\limits_{}^{} \frac{u \cdot 2udu}{u^2-4}\] \[=\int\limits_{}^{}\frac{2(u^2-4)+8}{u^2-4} du=\int\limits_{}^{}(2+\frac{8}{u^2-4}) du\] \[2u+8\int\limits_{}^{} \frac{1}{(u-2)(u+2)} du =2 \sqrt{4+e^x} +8\int\limits_{}^{}(\frac{A}{u-2}+\frac{B}{u+2}) du \] .... and blah blah is the way i would have done it my computer is freaking out so i can't finish

OpenStudy (jamesj):

Yes, we have exactly the same thing. Like I said, this was the last problem in my high school mathematics text book. I was very happy when I solved it the first time! :-)

myininaya (myininaya):

i mean is the samething james has basically